OpenStudy (anonymous):

Approximate the real number solutions to the polynomial function f(x) = x3 + 4x2 – 4x – 16 .

OpenStudy (anonymous):

at least 1

OpenStudy (anonymous):

at most 3

OpenStudy (anonymous):

huh ? @Jhannybean help ?

OpenStudy (jhannybean):

remember our last problem? the number of zeroes is determined by the variable with the highest power in a polynomial? What he's saying it you can have AT MOST 3 zeroes of a function (\(\large \color{green}{x^3}+4x^2-4x+16\)) and can have AT LEAST 1 after you break it down to the factored form.

OpenStudy (anonymous):

i suggest grouping

OpenStudy (anonymous):

but this questions is kind of different, thats why im confused.

OpenStudy (anonymous):

i dont get what to do.

OpenStudy (jhannybean):

@completeidiot is better at factoring this than i am :) haha. still learning.

OpenStudy (anonymous):

factor x^3+4x^2 also factor -4x-16

OpenStudy (anonymous):

-4(x+4) im confused.

OpenStudy (anonymous):

f(x)=x^3+4x^2-4x-16=x^2(x+4)-4(x+4)=(x+4)(x^2-4)=(x+4)(x-2)(x+2)=0 so x=-4, x=2, x=-2

OpenStudy (anonymous):

thankkyyoooouuuu !!!